# Biblio.bib

### v2

@@ 3.2. Stability estimate
In this subsection, we show that Theorem 2 follows from the microlocal approach if $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q$ is solenoidal injective.
@@
$\langle\pi_{m(g_{0})}^{*}Q(g-g_{0}),1\rangle_{L^{2}(SM)}\leq C_{n}|\mathrm{Vol}_{g}(M)-\mathrm{Vol}_{g_{0}}(M)|+C_{n}\|g-g_{0}\|_{C^{5,\alpha}}^{2}.$
@@
[…]
@@
Here, we used that $\nabla g_{0}=0$. Now, we follow the argument of [GuKnLef, Proposition 3.5] to obtain:
@@
$\langle\pi_{4}^{*}Q(S),1\rangle_{L^{2}(SM)}=\langle\pi_{4}^{*}Q(S),\pi_{4}^{*}L(g_{0})\rangle_{L^{2}(SM)}=C_{n}\langle Q(S),L(g_{0})\rangle_{L^{2}(S^{4}T^{*}M)},$
for some constant $C_{n}>0$ that we do not need to make explicit. We note that for any symmetric $2$-tensor $T$, a direct computation yields
@@
Here, we used the commutation relation between $\operatorname{tr}$ and its adjoint $L$ (see for instance [GuMaz, Equation $(4.19)$ Chapter $4.1$]) which gives
@@
where we used that $\langle S\circ J,g_{0}\rangle=\langle S,g_{0}\rangle$, did the same calculation as in the hyperbolic case and where $C_{n},C_{n}^{\prime}>0$ are some constants depending on $n$. In particular, this first part of $Q(S)$ can be bounded just like in the first case.
@@
we used the fact that $g_{0}\circ J=g_{0}$ as well as the fact that $L$ and $J$ commute. Now, we can use the commutation relation between $D_{g_{0}}^{*}$ and $L$ (see for instance [GuMaz, Corollary 4.13]) which gives, for some constant $C_{n}$,
@@
For the octonionic case, the scalar products we are yet to bound are $\langle L(R^{\circ}(S)),L(g_{0})\rangle$, $\langle\mathrm{Sym}(\nabla^{2}S),L(g_{0})\rangle$ as well as $\langle\mathrm{Sym}(R^{\circ}(\nabla^{2}S)),L(g_{0})\rangle$. For the first scalar product, we first note from (2.47) that $R^{\circ}$ is selfadjoint. Now,
@@
for some constant $C_{n}>0$. We use [Ruan, Corollary 5.5] to get
@@
From (2.7), this implies $R^{\circ}(L(g_{0}))=-36g_{0}$. This shows that this pairing vanishes and does not contribute to the average. For the second term, we use the fact that $\mathrm{Sym}$ is an orthogonal projection to obtain
@@
We will compute the principal symbol of $Q$ in each cases. Note that by [Lef, Lemma 14.1.9], $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}$ is a pseudo-differential operator of order 00 with principal symbol equal to $\sigma_{\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}}(x,\xi)=\pi_{\mathrm{Ker}(\iota_{\xi})}$ where $\iota_{\xi}$ denotes the contraction with the tangent vector $\xi^{\sharp}$ and $\pi_{\mathrm{Ker}(\iota_{\xi})}$ is the projection onto ${\mathrm{Ker}(\iota_{\xi})}$. We will thus obtain the principal symbol of $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q$ by composing with the principal symbol of $Q$.
@@
which proves that $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q$ is elliptic on solenoidal tensors, i.e its principal symbol satisfies $\sigma_{\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q}(x,\xi)\geq c\pi_{\mathrm{ker}(\iota_{\xi})}\|\xi\|^{2}$ for some $c>0$ and $\|\xi\|\gg 1$.
If $g_{0}$ is either complex hyperbolic or quaternionic hyperbolic, then we prove that $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q$ is of gradient type on $\mathrm{Ker}(D_{g_{0}}^{*})$, i.e its principal symbol is injective on $\mathrm{Ker}(D_{g_{0}}^{*})$ at all points. Using (3.3) or (3.4), we see that
@@
with $m=1$ for a quotient of a complex hyperbolic space and $m=3$ for a quotient of the quaternionic hyperbolic space. The principal symbol of $\nabla^{*}\nabla$ is $\|\xi\|^{2}$ and using [Lef, Lemma 14.1.8], the principal symbol of $D_{g_{0}}$ is $ij_{\xi}:=i\mathrm{Sym}(\xi^{\sharp}\cdot)$ where $\xi^{\sharp}$ is the image by the musical isomorphism of the co-vector $\xi$. In total, this gives
@@
see [Lef, Lemma 5.1.15]. Fix an orthonormal basis $(e_{i})_{i=1}^{n}$ such that $v_{1}=v$ and $v_{1},\ldots,v_{8}$ span $\mathrm{Cay}(v)$. Using (3.11),
$-\pi_{2}^{*}\mathrm{Sym}(R^{\circ}(\xi(\cdot)^{2}S))(x,v)=\sum_{i=1}^{n}\xi(R(e_{i},v)v)\xi(e_{i})S_{x}(v,v)$
@@
In all the previous cases, this means that $(\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q)^{*}\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q$ is elliptic on solenoidal tensors and thus there exists a sharp parametrix (see [Lef, Theorem 6.1.1]), i.e there exists $A\in\Psi^{-2}$ a pseudo-differential operator of order $-2$ such that
@@
Under the hypothesis of Theorem 2, if $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q$ is solenoidal injective ((((where $Q$ is defined in Theorem 6)))) then one has (1.3).
@@
Let $g\in\mathcal{U}$, we can use a Taylor expansion as well as the slice Lemma 2.3 to obtain for any $\alpha\in(0,1)$ a diffeomorphism $\phi$ such that
@@
Suppose by contradiction that there exists $S\in\mathrm{Ker}(D_{g_{0}}^{*})$ such that $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}QS=0$. In particular, one has
@@
The real dimension $n$ is even in this case. Note however that any quotient of $\mathbb{C}\mathbb{H}^{1}$ is actually isometric to a real-hyperbolic case, see [BriHa, Proposition 10.12]. In the following, we will thus suppose without loss of generality that $n\geq 4$.
@@
Strategy for the solenoidal injectivity in the complex case. The idea is to first consider $p\in C^{\infty}(M;S_{0}^{3}T^{*}M)$ the trace-free part of $\tilde{p}$ and project $Q(S)=D_{g_{0}}\tilde{p}$ onto the trace-free component. In particular, using the Pestov identity (2.12) will provide an expression of $p$ in terms of $S$. Now, this means that using the commutativity relation $[D_{g_{0}},L]=0$ :
@@
Solving for the trace-free part of $\tilde{p}$. We then use (2.28) to obtain $\pi_{4}^{*}(D_{g_{0}}^{J}D_{g_{0}}^{J}S)=-H^{2}\pi_{2}^{*}S.$ In particular, projecting the equation $Q(S)=D_{g_{0}}\tilde{p}$ onto $E_{4}^{4},E^{2}_{4}$ and $E^{0}_{4}$ yields, with the notations of Lemma 2.8,
@@
Computing $Q(S)-D_{g_{0}}p$ as a function of $S$. In particular, this means that one has, using $[D_{g_{0}},L]=0$, for some one-form $q$,
@@
[…]
(4.12) $\begin{split}\langle V(f,h),f\rangle&=\langle(H_{+}H_{-}+H_{-}H_{+})f+H_{+}H_{+}h+X_{-}X_{+}Jf+X_{-}H_{+}Vf,f\rangle\\ &=-\|Hf\|^{2}-\langle H_{+}h,H_{-}f\rangle-\langle X_{+}Jf,X_{+}f\rangle-\langle H_{+}Vf,X_{+}f\rangle\end{split}$
@@
In the previous computation, if we applied the second bound of Lemma 2.9 ((((instead of the first one)))) to the term $\|Xf\|_{L^{2}(SM)}^{2}+\|Hf\|_{L^{2}(SM)}^{2}$, we would obtain a lower bound of the form $-O\big{(}\tfrac{1}{n}\big{)}\|\nabla_{g_{0}}S_{0}\|_{L^{2}(S^{2}T^{*}M)}^{2}$ in (4.10). This would give an asymptotically better lower bound ((((when $n\to+\infty)$. Nevertheless, this would not make the following Weitzenböck argument work for small values of $n$ so we refrain from using it.
[…]

### v3

@@ 3.2. Stability estimate
In this subsection, we show that Theorem 2 follows from the microlocal approach if $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q$ is solenoidal injective. The solenoidal injectivity for real and complex hyperbolic metrics will be shown in Section 4.
@@
$\langle\pi_{4}^{*}Q(g-g_{0}),1\rangle_{L^{2}(SM)}\leq C_{n}|\mathrm{Vol}_{g}(M)-\mathrm{Vol}_{g_{0}}(M)|+C_{n}\|g-g_{0}\|_{C^{5,\alpha}}^{2}.$
@@
In the following proof, except if stated otherwise, the notation $\langle\cdot,\cdot\rangle$ will denote the scalar product induced on symmetric tensors. If $g_{0}$ is a hyperbolic metric, then from (3.2), one has
@@
Here, we used that $\nabla g_{0}=0$. Now, we follow the argument of [16, Proposition 3.5] to obtain:
@@
$\langle\pi_{4}^{*}Q(S),1\rangle_{L^{2}(SM)}=\langle\pi_{4}^{*}Q(S),\pi_{4}^{*}L(g_{0})\rangle_{L^{2}(SM)}=\Lambda_{4}^{n}\langle Q(S),L(g_{0})\rangle_{L^{2}(S^{4}T^{*}M)}.$
We will write $Q(S)=L(T(S))-\tfrac{1}{8}D_{g_{0}}^{J}D_{g_{0}}^{J}(S)$ for a symmetric $2$-tensor $T(S)$ defined in (3.3). We note that for any symmetric $2$-tensor $T$, a direct computation yields
@@
Here, we used the commutation relation between $\operatorname{tr}$ and its adjoint $L$ (see for instance [18, Equation $(4.19)$ Chapter $4.1$]) which gives
@@
where we used that $\langle S\circ J,g_{0}\rangle=\langle S,g_{0}\rangle$, did the same calculation as in the hyperbolic case and where $C_{n},C_{n}^{\prime}>0$ are some constants depending on $n$. In particular, the term $L(T(S))$ can be bounded just like in the first case.
@@
we used the fact that $g_{0}\circ J=g_{0}$ as well as the fact that $L$ and $J$ commute. Now, we can use the commutation relation between $D_{g_{0}}^{*}$ and $L$ (see for instance [18, Corollary 4.13]) which gives, for some constant $C_{n}$,
@@
For the octonionic case, the scalar products we are yet to bound are $\langle L(R^{\circ}(S)),L(g_{0})\rangle$, $\langle\mathrm{Sym}(R^{\circ}(\nabla^{2}S)),L(g_{0})\rangle$ as well as $\langle\mathrm{Sym}(\nabla^{2}S),L(g_{0})\rangle$. For the first scalar product, we first note from (2.48) that $R^{\circ}$ is selfadjoint. Now,
@@
for some constant $C_{n}>0$. We use [33, Corollary 5.5] to get
@@
From (2.7), this implies $R^{\circ}(L(g_{0}))=-36g_{0}$. This shows that this pairing vanishes and does not contribute to the average. For the last term, we use the fact that $\mathrm{Sym}$ is an orthogonal projection to obtain
@@
Note that up to here, using the computation of the Hessian (3.1), Lemma 3.2 and the coercive estimate (2.19) gives
(3.15) $\begin{split}\|\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q(S)\|_{H^{-1/2}}^{2}\leq C_{n}\langle d^{2}\Phi(g_{0})S,S\rangle&+C_{n}\big(|\mathrm{Vol}_{g}(M)-\mathrm{Vol}_{g_{0}}(M)|\\ &+\|g-g_{0}\|_{C^{5,\alpha}}^{2}\big)^{2}.\end{split}$
@@
We will compute the principal symbol of $Q$ in each case. Note that by [27, Lemma 14.1.11], $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}$ is a pseudo-differential operator of order 0 with principal symbol equal to $\sigma_{\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}}(x,\xi)=\pi_{\mathrm{Ker}(\iota_{\xi})}$ where $\iota_{\xi}$ denotes the contraction with the tangent vector $\xi^{\sharp}$ and $\pi_{\mathrm{Ker}(\iota_{\xi})}$ is the projection onto $\mathrm{Ker}(\iota_{\xi})$}. We will thus obtain the principal symbol of $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q$ by composing with the principal symbol of $Q$.
@@
which proves that $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q$ is elliptic on solenoidal tensors, i.e., its principal symbol satisfies $\sigma_{\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q}(x,\xi)\geq c\pi_{\mathrm{ker}(\iota_{\xi})}\|\xi\|^{2}$ for some $c>0$ and $\|\xi\|\gg 1$.
If $g_{0}$ is either complex hyperbolic or quaternionic hyperbolic, then we prove that $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q$ is of gradient type on $\mathrm{Ker}(D_{g_{0}}^{*})$, i.e., its principal symbol is injective on $\mathrm{Ker}(D_{g_{0}}^{*})$ at all points. Using (3.3) or (3.4), we see that
@@
with $m=1$ for a quotient of a complex hyperbolic space and $m=3$ for a quotient of the quaternionic hyperbolic space. The principal symbol of $\nabla^{*}\nabla$ is $\|\xi\|^{2}$ and using [27, Lemma 14.1.11], the principal symbol of $D_{g_{0}}$ is $ij_{\xi}:=i\mathrm{Sym}(\xi^{\sharp}\cdot)$ where $\xi^{\sharp}$ is the image by the musical isomorphism of the co-vector $\xi$. In total, this gives
@@
see [27, Lemma 5.1.15]. Fix an orthonormal basis $(e_{i})_{i=1}^{n}$ such that $v_{1}=v$ and $v_{1},\ldots,v_{8}$ span $\mathrm{Cay}(v)$. Using (3.12),
$\pi_{2}^{*}\mathrm{Sym}(R^{\circ}(\xi(\cdot)^{2}S))(x,v)=\sum_{i=1}^{n}\xi(R(e_{i},v)v)\xi(e_{i})S_{x}(v,v)$
@@
In all the previous cases, this means that $(\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q)^{*}\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q$ is elliptic on solenoidal tensors and thus there exists a sharp parametrix (see [27, Theorem 6.1.3]), i.e., there exists $A\in\Psi^{-2}$ a pseudo-differential operator of order 2-2 such that
@@
Under the hypothesis of Theorem 2, if $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q$ is solenoidal injective ((where $Q$ is defined in Theorem 5)) then one has (1.3).
@@
Let $g\in\mathcal{U}$. We can use a Taylor expansion as well as the slice Lemma 2.3 to obtain for any $\alpha\in(0,1)$ a diffeomorphism $\phi$ such that
@@
Suppose that there exists $S\in\mathrm{Ker}(D_{g_{0}}^{*})$ such that $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}QS=0$. In particular, one has
@@
The real dimension $n$ is even in this case. Note however that any quotient of $\mathbb{C}\mathbb{H}^{1}$ is actually isometric to a real-hyperbolic case, see [3, Proposition 10.12]. In the following, we will thus suppose without loss of generality that $n\geq 4$.
@@
Strategy for the solenoidal injectivity in the complex case. The idea is to first consider $p\in C^{\infty}(M;S_{0}^{3}T^{*}M)$ the trace-free part of $\tilde{p}$ and project $Q(S)=D_{g_{0}}\tilde{p}$ onto the trace-free component. In particular, using the Pestov identity (2.12) will provide an expression of $p$ in terms of $S$. Now, this means that using the commutativity relation $D_{g_{0}},L]=0$ :
@@
Solving for the trace-free part of $\tilde{p}$. We will now use the theory of raising/lowering operators defined in Subsection 2.4.
Using (2.10) and (2.35), we obtain $\pi_{4}^{*}(D_{g_{0}}^{J}D_{g_{0}}^{J}S)=-H^{2}\pi_{2}^{*}S.$ As a consequence of the mapping properties of $X$ recalled in Subsubsection 2.1.3, the only contribution for $\Omega_{4}$ are $X_{+}p$ for $\pi_{4}^{*}D_{g_{0}}\tilde{p}$ and $\tfrac{1}{8}H_{+}H_{+}f$ for $\pi_{4}^{*}Q(S).$
Next, using the notations of Lemma 2.7, the space $\Omega_{4}$ further splits into
$\Omega_{4}=E_{4}^{4}\oplus E_{4}^{2}\oplus E_{4}^{0}\oplus E_{4}^{-2}\oplus E_{4}^{-4}.$
In particular, projecting the equation $Q(S)=D_{g_{0}}\tilde{p}$ onto $E_{4}^{4},E^{2}_{4}$ and $E^{0}_{4}$ yields, using the mapping properties proved in Lemma 2.7,
@@
Computing $Q(S)-D_{g_{0}}p$ as a function of $S$. In particular, this means that one has, using $D_{g_{0}},L]=0$, for some one-form $q$,
@@
This allows us to do the calculations using spherical harmonics and in the following computations, $\langle\cdot,\cdot\rangle$ will denote the scalar product on $L^{2}(SM)$. We compute the first term using (4.3), the fact that $X_{+}^{*}=-X_{-}$ and $H_{+}^{*}=-H_{-}$,
(4.12) $\begin{split}\langle V(f,h),f\rangle&=\langle(H_{+}H_{-}+H_{-}H_{+})f+H_{+}H_{+}h+X_{-}X_{+}Jf+X_{-}H_{+}Vf,f\rangle\\ &=-\|H_{+}f\|^{2}-\langle X_{+}Jf+H_{+}Vf,X_{+}f\rangle-\|H_{-}f\|^{2}-\langle H_{+}h,H_{-}f\rangle\end{split}$
@@
In the previous computation, if we applied the second bound of Lemma 2.8 ((instead of the first one)) to the term $\|Xf\|_{L^{2}(SM)}^{2}+\|Hf\|_{L^{2}(SM)}^{2}$, we would obtain a lower bound of the form $O\big(\tfrac{1}{n}\big)\|\nabla_{g_{0}}S_{0}\|_{L^{2}(S^{2}T^{*}M)}^{2}$ in (4.10). This would give an asymptotically better lower bound ((when $n\to+\infty)$. Nevertheless, this would not make the following Weitzenböck argument work for small values of $n$ so we refrain from using it.
Using a Weitzenböck formula. We conclude the proof of Proposition 4.2. In the following, $\langle\cdot,\cdot\rangle$ denote the scalar product on $L^{2}(S^{2}T^{*}M)$. We use the fact that $W(S)=Q_{1}(S)+\tfrac{1}{8}Q_{0}(S)$, equation (4.9) and the lower bound (4.10):